OpenStudy (anonymous):

What is the relationship between the lines determined by the following two equations? 15x – 3y = –12 y = 5x + 7 A. parallel B. perpendicular C. they are the same line D. neither parallel nor perpendicular

OpenStudy (anonymous):

parallel bcoz their slopes are same

OpenStudy (zale101):

since the second equation is already in this y = mx + b, y = 5x + 7 we already know that the slope is 5 15x-3y=-12 Change to 3y=15x-12 in slope form so divide 15x-12 by 3 you’ll get y= 15x/3-12/3 then simplify,, you’ll get y=5x-4 the slope is also 5 so both equations have the same slope

OpenStudy (anonymous):

First one is easier to understand :P